# Comparative Study of State-based Neural Networks for Virtual Analog Audio Effects Modeling

### v5

@@ 4 Results
Figure 4: Comparison of the RMSE between the target and predicted output. The RMSE is computed using windows of $4096$ samples with $75\%$ overlap. When the model includes conditioning, the results refer to the highest conditioning values, representing the most challenging scenario.
@@
In the OD300 and Neutron’s overdrive dataset, the LSTM model presents the lowest MSE and ESR, indicating better general amplitude matching between target and prediction. On the other hand, the S4D and S6 present better accuracy when considering the transients, the RMS envelope, and frequency-related content. This can be observed from the general lower error in terms of RMSE, spectral flux, STFT, and MFCC. In the case of maximum distortion, we can notice LRU prediction having significantly lower energy than other models. Notably, the spectrograms shown in Figure 5 illustrate that S4D and S6 models better match the energy in the harmonics at higher frequency regions. LRU fails to generate the harmonics in the OD300 case while ED presents the opposite behavior in Neutron’s overdrive. Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the investigated architectures. This challenge can also be seen in the figures. The overdrive appears to be cutting the higher frequencies, behaving also as a low-pass filter.
@@
For models trained on the Saturator dataset, the LRU model presents lower MSE and $M_{RMSE}$ values, but when considering other metrics, no models appear clearly outperforming. When inspecting Figures  4, LRU is again left behind, while other models present closer performances. In contrast, LRU achieves closer matching in the spectrograms, visible in Figure 5. Other models tend to generate more harmonic, especially LSTM that introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance.
When training models with the Equalizer dataset, LSTM performs better in the time domain, having lower MSE and ESR errors while performing poorly in the frequency domain. Figures  4 show close performance among the models, while looking at the spectrograms S6 present better matching. The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters. This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling. Among the tested models, the LSTM and LRU are the least accurate in the time domain metrics, exhibiting higher loss in both the LA-2A and CL 1B cases. In contrast, in the frequency domain, the gap is less prominent. The S6 and ED models demonstrate superior accuracy for these two tasks. Specifically, S6 outperforms the ED in the LA-2A case, while the ED excels in the CL 1B case. S4D presents similar performance than S6.
@@
In the LA-2A case, we observe an overall good match for RMSE in Figure 4. The LRU model displays slightly more mismatches in both plots. Upon examining Figure  5, spectrograms related to the S6 model have the best match with the target output, while the other models present similar mismatches when looking at the higher frequencies. More difficult is the CL 1B case, where LRU presents the best envelope matching in Figure 4, and Figure  5. In the case of the compressor, the capability to encode past information of the signal is critical, LSTM is confirmed to have a worse temporal tracking ability than the other models, which present design improvements. These improvements are particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters. These reach up to $300$ ms and $5$ seconds, respectively.
Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal. Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 5. However, when considering the quantitative metrics, the LRU model outperforms the others. The LSTM model, in particular, tends to incorrectly predict excessive energy on the RMS envelope, as visible in Figure 4, and to introduce artifact, as shown in Figure 5.
Generally, performance metrics are relatively close across models across most effects examined in this study. Although the S4D model slightly outperforms other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study. The S6 model often performs very closely, while LRU presents the bigger variance across the different metrics and datasets.
The ED shows its quality when modeling the compressors. The ED, LRU, and particularly the S4D and S6 models are more effective at encoding past signal information within their internal states. This is suggested by the case of the cases of optical compressors, where they significantly outperform other models. Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters. Moreover, LSTM models tend to introduce artifacts in some cases.
@@ 5 Conclusion
The computational challenges of simulating the nonlinearities of analog effect, described by complex differential equations, are a significant obstacle for real-time applications. The complexity of these nonlinearities makes them difficult to model accurately, but they also contribute to the unique sound characteristics of the devices, making them appealing for both academic and industrial purposes. Modeling these nonlinear aspects is a difficult task that requires simplifications and assumptions to ensure stability and efficiency in the numerical schemes. Artificial neural networks (ANN) have emerged as a popular data-driven approach for this modeling process. While they offer a convenient and automated approach, they may struggle with the computational complexity and flexibility required for handling control parameters. In some cases, hybrid solutions that incorporate physics knowledge or measurements from internal components may provide advantages. However, there are situations where black-box models are necessary and feasible. Recurrent neural networks (RNNs) utilize internal states as memory mechanisms, allowing them to encode the signal history. This feature enables more interactive solutions as their memory capacity is not solely dependent on the input amount. In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Space models (SSM), which have demonstrated remarkable results in sequence modeling. To carry out this study, we design five models incorporating parameter conditioning and providing low-latency response. The control parameters are integrated into the models using the Feature-wise Linear Modulation (FiLM) method. Our selection of effects includes overdrive, saturation, low pass filter, equalization, and compressors. We evaluate these models using various metrics, which provide valuable insights for future research and development. Specifically, our metrics consider the root-mean-squared energy envelopes, the transients, and the frequency content.
LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively. Conversely, LSTMs tend to introduce more artifacts than other models considered in this study. In contrast, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression. In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history. Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced. Lastly, LRUs are less consistent in their quality across the different metrics.
In this study, we did not carry out architecture—and effect-specific optimizations, which can definitely boost the accuracy, such as tuning hyperparameters and changing the size of the various layers. We limited computational complexity and input-output latency and focused the study on models that can be theoretically implemented in consumer-grade devices for live audio applications.

### v6

@@ 4 Results
Figure 4: Comparison of the RMS energy between the target and predicted output. The RMS is computed using windows of $4096$ samples with $75\%$ overlap. When the model includes conditioning, the results refer to the highest conditioning values, representing the most challenging scenario.
In the OD300 and Neutron’s overdrive dataset, the LSTM model demonstrates superior overall performance. However, it does not achieve the lowest $M_{STFT}$, although its values are close to the best. Additionally, S6 struggles more with emulating both effects, particularly with the Neutron’s overdrive dataset. When trained on the Saturator dataset, S4D, S6, and LRU deliver the strongest performance, with LSTM and ED slightly trailing. In contrast, for the Equalizer dataset, LSTM consistently performs better than the other models.
The optical compressor features extended transients and lengthy time dependencies between the input signal and the alteration applied by the device, influenced by its adjustable attack and release parameters. This complexity is especially pronounced with the CL 1B compressor, presenting a challenging task for ANN modeling. Of the tested models, LSTM is the least accurate, with higher losses in both the LA-2A and CL 1B scenarios. In contrast, ED, S4D, and S6 demonstrate superior accuracy for these tasks, with S6 excelling in the LA-2A case and ED in the CL 1B case.
In analyzing the metrics of models trained to emulate the low-pass filter, both S4D and ED show the best performance. However, all models reveal a considerable frequency mismatch with the target output. This suggests their inability to model the signal alteration process of this parametric effect, which selectively removes high-frequency components while leaving those below the cutoff frequency unchanged.
Overdrive effects significantly distort the input signal, leading to the introduction of harmonic components. In these cases, LSTM-based models have been shown to be more effective. Conversely, the saturator determines mild distortion, similar to tape processing, by smoothing out the transients of the input signal. This situation may involve hysteresis and short delays associated with magnetic recording  Mikkonen et al., (2023), leading to longer temporal dependencies between input and output compared to distortion effects. This characteristic may explain the advantages observed with SSM-based layers. However, we lack information about the specific implementation of the software used, preventing us from confirming this hypothesis. Conversely, The Equalizer introduces tube-like saturation, which may explain the better performance of LSTMs in this context. SSM-based layers also demonstrated improved performance with optical compressors, where the effective management of transients and temporal dependencies is crucial.
When ranging the models according to performance metrics, inconsistencies arise with $M_{SF}$ and $M_{STFT}$, as these metrics don’t always align with the trends observed in time-based metrics. Nonetheless, no distinct patterns are apparent. This could be due to the loss function’s emphasis on the time domain, where variations in values are typically subtle.
To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the five distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMS energy in Figure  4, and the spectrograms in Figure  5. For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $10$ second input signal containing bass, guitar, and drum loop sound. These input signals are part of the respective test sets.
@@
In the case of maximum distortion, we can notice LRU prediction having significantly lower energy than other models. Notably, the spectrograms shown in Figure 5 illustrate that S4D and S6 models better match the energy in the harmonics at higher frequency regions. LRU fails to generate the harmonics in the OD300 case, while ED presents the opposite behavior in Neutron’s overdrive. Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the investigated architectures. This challenge can also be seen in the figures. The overdrive appears to be cutting the higher frequencies, behaving also as a low-pass filter.
@@
When inspecting models emulating the Saturator, LRU is again left behind, while other models present closer performances. In contrast, LRU achieves closer matching in the spectrograms, visible in Figure 5. Other models tend to produce more harmonics, particularly the LSTM, which introduces noticeable distortions that are clearly audible in the audio output and adversely impact its performance. Models emulating the Equalizer demonstrate similar performance across different designs. However, when examining the spectrograms, S6 shows superior matching to the target audio characteristics.
@@
In the LA-2A case, we observe an overall good match for RMSE in Figure 4. The LRU model displays slightly more mismatches in both plots. Upon examining Figure  5, spectrograms related to the S6 model have the best match with the target output, while the other models present similar mismatches when looking at the higher frequencies. More difficult is the CL 1B case, where LRU presents the best energy matching in Figure 4, and Figure  5. In the case of the compressor, the capability to encode past information of the signal is critical, and LSTM is confirmed to have a worse temporal tracking ability than the other models, which present design improvements. These improvements are particularly evident with the CL 1B when setting long attack and release times through the conditioning control parameters. These reach up to $300$ ms and $5$ seconds, respectively. In the case of the low-pass filter, input-output signal alterations are the most pronounced and diverse among the effects considered, challenging the network’s ability to learn this mapping. The difficulty mainly stems from the filter’s range, which varies from a straightforward pass-through at high cutoff frequencies to a highly selective filter at low cutoff frequencies. Nevertheless, informal listening of the predicted outputs indicates that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 5. The LSTM model, in particular, tends to incorrectly predict excessive energy on the RMS energy, as visible in Figure 4, and to introduce distortions, as shown in Figure 5.
Generally, performance metrics are relatively close across models across most effects examined in this study. When looking at objective metrics, no models appear clearly outperforming the others, but all share similar learning capability. MSE, ESR, and $M_{NRMSE}$ errors present the same trends, while performing better in the time domain does not lead to the same in the frequency domain. The S4D and S6 models often yield similar performance, while LRU exhibits the greatest variance across different metrics and datasets. When modeling compressors, S4D, S6, and ED demonstrate high quality. The ED, LRU, and especially S4D and S6 models excel at encoding historical signal information within their internal states, significantly outperforming LSTM models in compressor scenarios. Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical transients and time dependencies, such as the overdrives. Furthermore, in certain instances, LSTM models tend to introduce significant distortion.
Table 4 presents the $p$-values calculated using the Friedman test, which determines whether there are statistically significant differences between the medians of samples. All values are below $0.05$, indicating statistical significance in the results. The Wilcoxon results are omitted for brevity, as they involve computing $p$-values for each pair of models trained on the same dataset. However, we observe $p$-values exceeding $0.05$ solely when comparisons involve LRU or when models are trained with the low-pass filter dataset. This supports the previously mentioned difficulty in learning low-pass filters and the high variance in the LRU model’s performance.
Table 4: $p$-values calculated from the Friedman test for each metric and dataset presented in Table 3, indicating the statistical significance of the rankings.
Dataset MSE ESR $M_{NRMSE}$ $M_{SF}$ $M_{STFT}$
OD300 $0.001$ $0.001$ $0.0008$ $0.011$ $0.0008$
Neutron Overdrive $0.001$ $0.001$ $0.0004$ $0.0006$ $0.0010$
Helper Saturator $0.002$ $0.002$ $0.002$ $0.001$ $0.0004$
Pultec Passive EQ $0.002$ $0.002$ $0.001$ $0.0007$ $0.013$
LA-2A $0.002$ $0.002$ $0.012$ $0.0006$ $0.002$
CL 1B $0.002$ $0.002$ $0.002$ $0.001$ $0.0004$
Neutron Filter $0.002$ $0.002$ $0.004$ $0.018$ $0.002$
@@ 5 Conclusion
Modeling analog effects is challenging, requiring simplifications to maintain numerical stability and efficiency. ANNs have gained popularity for their convenient, automated approach, but they can struggle with computational complexity and flexibility for control parameters. While hybrid solutions incorporating physics can offer advantages, black-box models are sometimes necessary. Recurrent neural networks, utilizing internal states for memory, provide interactive solutions by encoding signal history, enhancing their modeling capability beyond input size limitations.
In this paper, we conducted a comparative analysis to assess the performance of LSTM models against recently proposed recurrent models, specifically LRU and SSMs, which have shown impressive results in sequence modeling. Five models were designed with parameter conditioning and low-latency response using the Feature-wise Linear Modulation method, focusing on effects like overdrive, saturation, low-pass filter, equalization, and compressors. Our evaluation employed metrics considering root-mean-squared energy, transients, and frequency content for valuable insights into future research. Experimental results indicate that LSTM-based networks perform well in emulating distortion and equalizer effects. However, they tend to introduce noticeable distortions compared to other models in this comparative study. Conversely, LSTM structured in an encoder-decoder configuration, and SSMs, perform better in emulating tape saturation and compression effects, where encoding historical signal information may significantly impact the accuracy. However, the performance gap is not always pronounced. Regarding the low-pass filter, no architecture successfully learns its audio processing characteristics. In general, MSE, ESR, and $M_{NRMSE}$ errors present the same trends, while performing better in the time domain does not lead to the same in the frequency domain. Lastly, LRU shows inconsistent performance across different datasets.